# Rajatsubhra Chakraborty Arkaprava Sinha  Dominick Reilly  Manish Kumar Govind Pu WangFrancois BremondSrijan Das UNC Charlotte  Inria  Université Côte d’Azur

### v1

@@ 4.4 Comparison to the state-of-the-art
We compare LLAVIDAL against the state-of-the-art (SOTA) in the performance on video description generation and ADLMCQ tasks involving action recognition and forecasting.
Video Description Generation. Table 4 shows the performance comparison of baseline LLVMs and LLAVIDAL on their video description capabilities on the Charades and TSU datasets. Video-level descriptions are obtained directly from the Charades dataset. For the TSU dataset, comprising lengthy videos, we segment each video into 1-minute clips and input them individually to the LLVMs for generating clip-level descriptions. Subsequently, we concatenate all clip-level descriptions and utilize GPT-3.5 turbo to summarize them into a video-level description, following the same instruction template utilized in our dense description pipeline for ADL-X. LLAVIDAL consistently surpasses SOTA and outperforms all models including, image captioners-summarizers pipelines which are trained on billions of images, across all 5 VideoChatGPT metrics. However, in the Mementos Evaluation, LLVM baselines exhibit superior performance over LLAVIDAL in the Smarthome domain. This discrepancy may be attributed to the loss of relevant information when generating video-level descriptions using GPT.
Table 5: ADLMCQ - Action Recognition
Method Charades Smarthome
VideoLlama [19] 33.0 27.4
VideoLlava [18] 44.4 54.0
VideoChatGPT [20] 56.0 40.8
ADL-X ChatGPT [20] 58.0 52.3
LLAVIDAL 59.0 58.8
Table 6: ADLMCQ - Action Forecasting
Method LEMMA TSU
VideoLlama [19] 20.8 15.6
VideoLlava [18] 32.2 20.2
VideoChatGPT [20] 35.7 25.0
ADL-X ChatGPT [20] 44.8 25.3
LLAVIDAL 52.6 27.0
ADLMCQ. Table 6 compares LLAVIDAL to SOTA LLVMs on the ADLMCQ-AR benchmark. LLAVIDAL achieves significant improvements, surpassing VideoChatGPT by +5.4% and +44.1% on the Charades and Smarthome datasets, respectively. Similarly, Table 6 demonstrates LLAVIDAL’s superiority on the ADLMCQ-AF benchmark. It outperforms VideoChatGPT by up to +47.3%, highlighting its exceptional capability in action forecasting tasks.
Figure 4 provides a visual comparison of LLAVIDAL against representative baselines on the ADL benchmarks. More visual samples are provided in the Appendix G.
5 Conclusion & Future Work
In this work, we present a framework for curating ADL datasets for instruction tuning LLVMs, thus introducing ADL-X. We introduce LLAVIDAL, an LLVM capable of integrating 3d poses and human-object interaction cues by projecting their language contextualized representations into the LLM embedding space. To assess LLVM performance in ADL scenarios, we propose the ADLMCQ benchmark. Results demonstrate that LLAVIDAL, when trained on ADL-X, surpasses other LLVM baselines in ADLMCQ tasks, indicating its efficacy in grasping intricate temporal relationships within ADL contexts. Future research will focus on expanding ADL-X by integrating additional curated ADL datasets and exploring modality progressive training strategies to effectively integrate both pose and object cues within LLAVIDAL.
Acknowledgements
This work is supported in part by the National Science Foundation (IIS-2245652). Additionally, this material is based upon research in part supported by the Chateaubriand Fellowship of the Office for Science & Technology of the Embassy of France in the United States. We thank the department of computer science of UNC Charlotte for providing credits to access GPT 3.5 Turbo. We are grateful to Ahmed Helmy for supplying the essential GPUs. We thank Da Ma and Ezequiel Zamora of Wake Forest School of Medicine for sharing their insights.
Appendix
@@ Appendix A Overview
  • Section C: PoseCLIP
  • Section D: Additional Dataset Details
  • Section E: Additional Implementation Details
  • Section F: Improving Actions: Pose Cues vs Object Cues
  • Section G: Additional Qualitative Evaluation
  • Section H: LLM Prompts Used
  • Section I: Limitations
  • Section J: Licensing and Intended Use
@@ Appendix B Related Work
In this section, we delve into the recent datasets proposed for instruction tuning of LLVMs. We also present the recent advancements in multimodal conversational models both with image captioners and video encoders which consist of an LLM at the final stage to leverage its generation and linguistic understanding capabilities.
Data: Existing video-centric instruction datasets, such as VideoChat[17], Valley[26], Video-ChatGPT [20], and TimeChat [21], have made significant strides in advancing general video understanding and dialogue. However, these datasets exhibit limitations that render them inadequate for training LLVMs to understand with ADL. The primary issues lie in the insufficient task coverage, brevity of video lengths, and lack of real-world complexity that characterize ADL. While the TimeIT dataset from TimeChat offers improved video duration and task diversity compared to its predecessors [59, 20, 17, 26], it still falls short of fully capturing the intricacies and extended temporal nature of many multi-step ADL tasks. Similarly, ActivityNet [25], despite being a large-scale benchmark with 203 activity classes, falls short in terms of its applicability to ADL. While ActivityNet boasts a diverse taxonomy, the selected activity classes are not tailored to the ADL domain. The dataset’s focus on general video understanding does not guarantee sufficient representation of the unique challenges posed by ADL, such as intricate object interactions, fine-grained actions, and long-term temporal dependencies. It is to be noted that previous approaches like VideoChatgpt [20], VideoLlava [18] derive their instruction dataset from ActivityNet. Webvid, which is now de-comissioned due to privacy issues introduced in [60], consists of 2.5 million video-text pairs scraped from the web. Although it is a large-scale dataset, the videos are not specifically focused on ADL. The dataset covers a broad range of topics and video types, which may not adequately capture the nuances and challenges specific to ADL scenarios.
@@
Image captioners + LLM. Advancements in the abilities of LLMs in contextual understanding and language generation has led to the rise of multimodal conversational models. These methods, typically employ foundation models to generate visual features from images and project them to a space compatible with the language models. Flamingo [14] uses vision-language resampler in conjunction with gated cross-attention while BLIP2 introduces Q-Former map image features to the LLM embedding space. MiniGPT4 [10] uses a simple linear projection layer. However, these model fall short of becoming conversational assistants due to the absence of human instruction feedback. To this end, mPLUG-OWL [13] first aligns visual and linguistic features by multimodal autoregressive pretraining. It then performs mutlimodal instruction tuning with LoRA [61], which facilitates responses to be natural and aligned with human instructions. IntructBLIP [46] and LLaVA [12] introduce large scale human instruction datasets that facilitate LLM finetuning. PaLI [15] and Qwen-VL [16] are capable of direct training of the LLMs during pretraining or supervised finetuning stages. However, it leads to a loss of generalizability of the natural language capabilities of the LLM. CogVLM [44], on the other hand, introduces separate layers into the Transformer Block of the LLM to process image features using an independent QKV matrix and Feed Forward Network for images.
Large Language Vision Models (LLVMs). Researchers have been rigorously investigating methods to understand videos and develop video-conversational models integrated with large language models (LLMs). Methods like Socratic Models [62] and VideoChat [17], use pretrained foundation vision encoders [63, 64] along with LLMs to adapt them for video tasks.
Among dialog based models, VideoChatCaptioner [65] summarizes a video based on conversations between ChatGPT [7] and a captioner like BLIP2 [45], while ChatVideo [66] uses task-specific foundation models to create a database of "tracklets". A database manager and ChatGPT [7] work to generate responses from user queries during inference. Some approaches [67, 68] divide each video into segments and either option descriptions for each segment to be shared directly with LLMs or encode each segment, concatenate the tokens and project them to the LLM space. Models like [17, 19, 21] leverage Query Transformer (Q-Former) [45] for effective feature encoding and alignment. VideoLLaMA [19] first uses a vision transformer with an image Q-Former to obtain frame-level representations and then a video Q-Former for temporal modelling. TimeChat [21], which can encode variable length videos, uses a timestamp-aware frame encoder with a Q-Former to infuse temporal information into the vision tokens and subsequently a sliding window Q-Former to condense the frame-level features for the Projection Layer.
Similar to  [19], VideoLLaVA [18] jointly trains on images and videos, however, it pre-aligns the visual modalities to language using LanguageBind [69] encoders. VideoChatGPT [20] leverages both temporal and spatial features if a video, obtained by average pooling the frame-level features spatially and temporally, respectively. In contrast to these models, LLAVIDAL  incorporates both 3D Pose and object cues into LLaVA type conversational models. The integration of these cues is instrumental for the effective interpretation of ADL videos.
Appendix C PoseCLIP
PoseCLIP is a Pose-Language Model (PoseLM) that aligns 3D poses to a language contextualized space. It consists of two components: a Pose Encoder and a Text Encoder. Here we use a Hypergraph Transformer, Hyperformer [48] as the Pose Encoder, due to its ability to learn representations, based on the human kinematics and its capability of efficient skeleton action recognition. The pose encoder $f_{p}$ first processes each pose sequence, $P_{i}\in\mathbb{R}^{T_{p}\times 3\times J}$, where each frame within the sequence $T_{p}$ comprises $J$ joints. The Pose encoder generates frame-level pose representations which are aggregated using a Temporal Pooling Layer to obtain sequence-level representations $z_{i}^{p}$. The Text Encoder $f_{t}$ is derived from the CLIP [2] text encoder which is frozen and computes a text representation $z_{i}^{t}$ for the corresponding action label($t_{i}$). The PoseCLIP undergoes a two-stage training scheme. In the first stage, the pose encoder is pretrained on NTURGB+D [39], for Skeleton Action Recognition. In the second stage, the pose and text embeddings are aligned by maximizing their corresponding cosine similarities. The loss is given by 4.
$z_{i}^{p}=\frac{1}{T_{p}}\sum f_{p}(P_{i}),\hskip 5.42003ptz_{i}^{t}=f_{t}(t_{i})$ (3)
$\mathcal{L}_{CE}(z_{i}^{p},z_{i}^{t})=-\sum_{i}\log\frac{\exp(sim(z_{i}^{p},z^{t}_{i})/\tau)}{\sum_{j}\exp(sim(z_{j}^{p},z_{j}^{t})/\tau)}$ (4)
where $\tau$ is a temperature parameter and $sim(x,y)$ denotes the cosine similarity between $x$ and $y$.
Appendix D Additional Dataset Details
@@
Question Types. We divide our QA in different questions so that our model understands human object interaction holistically, we lay emphasis on actions performed and the sequence of actions occurring in the video and likewise how objects are associated with the actions. We carefully design such questions relevant to the videos with GPT 3.5 Turbo. The questions encompasses actions happening, summarization, objects in the scene, color of the objects and questions related to the video. For Pose as QA and Object as QA we construct two more questions each, for object we add "What are the relevant objects in the scene?" and "What is the object in the trajectory [x1,y1,x2,y2]?", for Pose we add "What is the motion of the body and joints relative to the actions?" and "Which joints are moving in the video?".
@@
Figure 5: Overview of ADL-X. Top Left: Training and test data distribution; Top Middle: Wordcloud of Textual Representation of Training Data; Bottom Left: Sample video frames with detected relevant object Bottom Middle: 3D Poses of the corresponding sample video; Right: Sample QA pairs
@@
Average video and sentence length. There is an average of $23$ words per sentence in our QA and average word count for each answer is $42$. The average video length is $10$ seconds in our dataset. We have $1262229$ nouns, $551172$ verbs, $40415$ actions and $722807$ objects in our QA showing the overall dynamics of dataset which is illustrated in WordCloud of the Figure 5.
Importance of cropping. When person-centric cropping is not applied to the videos in an ADL dataset, the resulting dense-level captions often include a significant amount of irrelevant information about the background scene. This extraneous information is not directly connected to the subject or the actions being performed, and its presence can introduce noise into the training data. If left unchecked, this noise can have a detrimental effect on the learning process, as the model may erroneously focus on the background details rather than the key elements of the ADL. By failing to isolate the relevant information, the model’s attention is diverted away from the crucial aspects of the task at hand, namely the individual performing the actions, the actions themselves, and the interactions between the person and objects in the scene. This dilution of focus can lead to suboptimal performance and hinder the model’s ability to accurately understand and classify ADLs. In contrast, by employing person-centric cropping, the irrelevant background information is effectively eliminated from the videos. This targeted approach ensures that the dense-level captions concentrate solely on the elements that are directly related to the subject and their actions. By maintaining this persistent focus on the relevant information, the training data becomes more coherent and informative, enabling the model to better capture the essential characteristics of the ADLs. In Fig 6, we illustrate an example why person centric cropping is important.
Appendix E Additional Implementation Details
We deployed a 4-bit quantized version of CogVLM-17B [44] for annotating frame-level captions. On an A5000 GPU, the inference uses 11GB of memory.The two prompts that are used to get the frame-level descriptions for the ADL-X are – "Give a detailed description of the actions happening and describe the image, include motions and the objects interacted by the person" and "Summarize the content of the image in details explaining all events happening". CogVLM uses Vicuna v1.5 7b [47] as their large language model and EVA2-CLIP-E [70] as their VIT encoder, the input image dimensions are $224\times 224$, the average time to annotate a video is 80 seconds at 0.5fps.
@@ Appendix E Additional Implementation Details
Figure 6: Left: uncropped videos and frame level annotations from CogVLM; Right: person centric cropping and CogVLM captions. The irrelevant information (marked red) adds noise to the annotations.
@@
LLAVIDAL details. To generate object cues, we perform frame-level object detection using BLIP2 and localization using OWLv2. BLIP2 [45] uses a ViT-L and a FlanT5 [71] architecture for detection, while OWLv2 [50] uses an OWL-ViT-L which is a CLIP based model for extracting localization features of the detected objects. In case of PoseCLIP, the Pose Encoder, Hyperformer, is pretrained on NTURGBD for 140 epochs for action recognition, and then is aligned with the CLIP Text Encoder for an additional 100 epochs. LLAVIDAL uses a Vicuna-v1.1 (7B) as the LLM which is frozen during instruction tuning.
Appendix F Improving Actions: Pose Cues vs Object Cues

### v3

@@ 7.5 Comparison to the state-of-the-art
We compare LLAVIDAL against the state-of-the-art on ADL MCQ tasks and ADL video description tasks.
ADL MCQ. Table 4 compares LLAVIDAL to state-of-the-art LLVMs on the ADL MCQ-AR and ADL MCQ-TC benchmarks. LLAVIDAL achieves substantial improvements in action recognition, outperforming the representative baseline, VideoChatGPT, by +8.2% and +21.4% on the Charades and Smarthome datasets, respectively. On the TSU dataset, LLAVIDAL surpasses VideoChatGPT by up to +82.7%, highlighting its superior temporal understanding. Figure 5 provides a visual comparison of LLAVIDAL against representative baselines on the ADL benchmarks, with additional examples in Appendix F.
ADL Video Description. Table 4 shows the performance comparison of baseline LLVMs and LLAVIDAL on their video description capabilities on the Charades and TSU datasets. Video-level descriptions are obtained directly from the Charades dataset. For the TSU dataset, comprising lengthy videos, we segment each video into 1-minute clips and input them individually to the LLVMs for generating clip-level descriptions. Subsequently, we concatenate all clip-level descriptions and utilize GPT-3.5 turbo to summarize them into a video-level description, following the same instruction template utilized in our dense description pipeline for ADL-X. Across all the 5 VideoChatGPT metrics, the results show that LLAVIDAL consistently surpasses current SOTA methods, and outperforms all models including the image captioners-summarizers pipelines which are trained on billions of images.
8 Conclusion
In this work, we present a novel semi-automated framework for curating ADL video instruction tuning dataset and introduce ADL-X. We introduce LLAVIDAL, an LLVM capable of integrating 3D skeletons and HOI modalities using a novel Multimodal Progressive training strategy. To assess LLVMs’ performance in ADL scenarios, we propose the ADL MCQ and video description benchmarks. Results demonstrate that LLAVIDAL, when trained on ADL-X, surpasses other baselines, indicating its efficacy in grasping intricate temporal relationships within ADL contexts.
Acknowledgments
This work is supported in part by the National Science Foundation (IIS-2245652) and NSF’s NAIRR Pilot initiative (NAIRR240338), which provided credits to access GPT-3.5 Turbo. This work was also supported by funds provided by The University of North Carolina at Charlotte.
Supplementary Material
@@ Appendix A Overview
  • Section C: Additional Dataset Details
  • Section D: Additional Implementation Details
  • Section E: Improving Actions: Skeleton Cues vs Object Cues
  • Section F: Additional Qualitative Evaluation
  • Section G: LLM Prompts Used
  • Section H: Limitations
  • Section I: Licensing and Intended Use
@@ Appendix B Related Work
In this section, we discuss the recent datasets proposed for instruction tuning of LLVMs. We also present the recent advancements in multimodal conversational models both with training-free methods and methods using visual connectors.
Table 5: Video Instruction Dataset Comparison.
Dataset Modalities Subjects Multiple Videos QA Pairs Atomic Actions Temporal Object Type Views per Vid Rand. Traj. TimeIT[56] RGB+L NA No 173000 173K Medium No No Web VideoChat[35] RGB+L NA No 8196 11K Low No No Web Valley[44] RGB+L NA No 64,687 65K Low No No Web VideoChatGPT [45] RGB+L NA No 27,801 100K Medium No No Web ADL-X RGB+S+L 106 Yes 16,343 100K High Yes Yes ADL
Data: Existing video-instruction datasets, such as VideoChat[35], Valley[44], Video-ChatGPT [45], and TimeIT [56], have made significant strides in advancing general video understanding and dialogue. Valley is derived from a website called Jukinmedia that provides videos with diverse categories and wide detailed descriptions. TimeIT dataset from TimeChat offers videos with temporal variations and task diversity. On the other hand, ActivityNet [8] boasts a diverse taxonomy with 203 activity classes, most activity classes are not tailored to the ADL domain. It is to be noted that most LLVMs like VideoChatgpt [45], VideoLlava [38] derive their instruction dataset from ActivityNet. Webvid, which is now de-comissioned due to privacy issues introduced in [3], consists of 2.5 million video-text pairs scraped from the web. Although these video instruction datasets are large in scale, offer diverse action classes, and focus on general video understanding, they fail to address the challenges specific to ADL. These challenges include intra-class temporal variations, long-term temporal relationships, complex human-object interactions, and videos captured in multiview settings. Unlike ADL videos, the internet videos in these datasets are predominantly shot by a cameraman, ensuring human-centered frames. Consequently, they do not capture the unstructured randomness spatially and temporally inherent in real-world ADL videos. In contrast, ADL-X is specifically designed to address the challenges inherent in ADL (see Table 5). It captures temporal unstructuredness through the temporal stitching of several unrelated actions in sequence and incorporates complex human-object interactions from the NTU120 dataset. Additionally, our proposed PAG and WS video description techniques effectively eliminate hallucinations, resulting in high-quality video-instruction pairs.
Image captioners + LLM. Advancements in Large Language Models (LLMs) have naturally extended vision-language models [52], leveraging LLMs to enhance reasoning capabilities. The emergence of these foundation models has given rise to training-free methods like Socratic Models [75] and VideoChat [35], which use pretrained vision encoders [34, 70] to map visual information into a language embedding space, followed by LLMs for downstream video tasks.
On the other hand, effective image captioners like CogVLM [68] introduce separate layers into the Transformer block of the LLM to process image features using independent QKV matrices and Feed Forward Networks specifically designed for images. Such effective captioning approaches have inspired methods that map visual information to language via image captioners, followed by processing with LLMs. Among dialog-based models, VideoChatCaptioner [10] summarizes videos through conversations between ChatGPT [7] and a captioner such as BLIP2 [33]. Similarly, ChatVideo [67] employs task-specific foundation models to create a database of ”tracklets,” with a database manager and ChatGPT [7] collaborating to generate responses for user queries during inference.
For long video understanding, approaches like [77, 71] segment videos into smaller units, providing either segment-level descriptions directly to LLMs or encoding each segment, concatenating tokens, and projecting them into the LLM space. Likewise, Language Repository [29] introduces write-and-read operations to prune text redundancies and extract information across various temporal scales. The Multimodal Video Understanding Framework [53] explores integrating video-specific information into an LLM-based framework by using off-the-shelf vision tools to extract three object-centric modalities from videos and fusing this information through natural language. Additionally, [51] investigates optimal strategies for key-frame selection to significantly reduce redundancies. However, despite these advances, training-free models fail to capture the complex temporal relationships intrinsic to ADL. These relationships, including long-term dependencies and intricate human-object interactions, remain a significant challenge for these approaches.
@@
Large Language Vision Models (LLVMs). The abilities of LLMs in contextual understanding and language generation have led to the introduction of video conversational models. These methods typically employ foundation models to extract visual features from images and project them into an embedding space compatible with language models. Flamingo [1] utilizes a vision-language resampler combined with gated cross-attention, while BLIP2 [33] introduces Q-Former to map image features into the LLM embedding space. Similarly, MiniGPT4 [81] uses a simple linear projection layer. However, these models fall short of becoming conversational assistants due to the lack of human instruction feedback. To address this, mPLUG-OWL [74] first aligns visual and language features through multimodal autoregressive pretraining, followed by multimodal instruction tuning using LoRA [25], enabling more natural and human-aligned responses. Models such as InstructBLIP [17] and LLaVA [40] introduce large-scale human instruction datasets to facilitate LLM fine-tuning. Meanwhile, models like PaLI [12] and Qwen-VL [2] allow direct training of LLMs during pretraining or supervised fine-tuning stages.
Other models, including VideoChat, VideoLLaMA, and TimeChat [35, 78, 56], leverage Q-Former for effective feature encoding and alignment. For example, VideoLLaMA [78] employs a vision transformer with an image Q-Former to obtain frame-level representations, followed by a video Q-Former for temporal modeling. Similarly, TimeChat [56] encodes variable-length videos using a timestamp-aware frame encoder with a Q-Former to infuse temporal information into vision tokens, followed by a sliding window Q-Former to condense frame-level features for the projection layer. Building on these approaches, VideoLLaVA [38] jointly trains on both images and videos, pre-aligning visual modalities to language using LanguageBind [80] encoders. VideoChatGPT [45] leverages both temporal and spatial features from videos, obtained by average pooling frame-level features both spatially and temporally.
In contrast to these models, LLAVIDAL incorporates 3D skeleton data and human-object interaction (HOI) cues alongside videos into the LLM embedding space. This integration enables the additional cues to learn discriminative video representations, making LLAVIDAL particularly effective for interpreting ADL videos, where temporal relationships and complex HOIs are crucial.
Figure 6: Overview of pipeline for Skeleton as QA and Skeleton as context.
Figure 7: Overview of pipeline for HOI as QA and HOI as context.
Appendix C Additional Dataset Details
@@
Question Types. We divide our QA in different questions so that our model understands human object interaction holistically, we lay emphasis on actions performed and the sequence of actions occurring in the video and likewise how objects are associated with the actions. We carefully design such questions relevant to the videos with GPT 3.5 Turbo. The questions encompasses actions happening, summarization, objects in the scene, color of the objects and questions related to the video. For Skeleton as QA and Object as QA, we construct two additional questions for each. For Skeleton, we include ”What is the motion of the body and joints relative to the actions?” and ”Which joints are moving in the video?”. For Object, we add ”What are the relevant objects in the scene?” and ”What is the object in the trajectory [x1, y1, x2, y2]?”. These are illustrated in Figures 6 and 7.
@@
Figure 8: Overview of ADL-X. Top Left: Training and test data distribution; Top Middle: Wordcloud of Textual Representation of Training Data; Bottom Left: Sample video frames with detected relevant object Bottom Middle: 3D skeletons of the corresponding sample video; Right: Sample QA pairs
@@
Average video and sentence length. There is an average of $23$ words per sentence in our QA and average word count for each answer is $42$. The average video length is $10$ seconds in our dataset. We have $126,2229$ nouns, $551,172$ verbs, $40,415$ actions and $722,807$ objects in our QA showing the overall dynamics of dataset which is illustrated in WordCloud of the Figure 8.
@@
Figure 9: Left: uncropped videos and frame level annotations from CogVLM; Right: PAG and CogVLM captions. The irrelevant information (marked red) adds noise to the annotations.
Importance of Person Augmented Generation (PAG). When PAG is not applied to the videos in an ADL dataset, the resulting dense-level captions often include a significant amount of irrelevant information about the background scene. This extraneous information is not directly connected to the subject or the actions being performed, and its presence can introduce noise into the training data. If left unchecked, this noise can have a detrimental effect on the learning process, as the model may erroneously focus on the background details rather than the key elements of the ADL. By failing to isolate the relevant information, the model’s attention is diverted away from the crucial aspects of the task at hand, namely the individual performing the actions, the actions themselves, and the interactions between the person and objects in the scene. This dilution of focus can lead to suboptimal performance and hinder the model’s ability to accurately understand and classify ADLs. In contrast, by employing person-centric cropping, the irrelevant background information is effectively eliminated from the videos. This targeted approach ensures that the dense-level captions concentrate solely on the elements that are directly related to the subject and their actions. By maintaining this persistent focus on the relevant information, the training data becomes more coherent and informative, enabling the model to better capture the essential characteristics of the ADLs. In Fig 9, we illustrate an example to highlight the importance of PAG in our semi-automated data curation framework.
Appendix D Additional Implementation Details
We deployed a 4-bit quantized version of CogVLM-17B [68] for annotating frame-level captions. On an A5000 GPU, the inference uses 11GB of memory.The two prompts that are used to get the frame-level descriptions for the ADL-X are – ”Give a detailed description of the actions happening and describe the image, include motions and the objects interacted by the person” and ”Summarize the content of the image in details explaining all events happening”. CogVLM uses Vicuna v1.5 7b [13] as their large language model and EVA2-CLIP-E [64] as their VIT encoder, the input image dimensions are $224\times 224$, the average time to annotate a video is 80 seconds at 0.5fps.
@@ Appendix D Additional Implementation Details
LLAVIDAL details. To generate HOI cues, we perform frame-level object detection using BLIP2 and localization using OWLv2. BLIP2 [33] uses a ViT-L and a FlanT5 [15] architecture for detection, while OWLv2 [47] uses an OWL-ViT-L which is a CLIP based model for extracting localization features of the detected objects. In case of skeletonCLIP, the skeleton Encoder, Hyperformer, is pretrained on NTURGBD for 140 epochs for action recognition, and then is aligned with the CLIP Text Encoder for an additional 100 epochs. LLAVIDAL uses a Vicuna-v1.1 (7B) as the LLM which is frozen during instruction tuning.
Appendix E Improving Actions: Skeleton vs HOI Cues
In this section, we present a comprehensive analysis of our multi-modal approach leveraging HOI tokens, skeleton tokens, and their progressive integration (MMPro) for fine-grained action recognition. Our empirical evaluation across a diverse set of household actions demonstrates distinct performance patterns across modalities, revealing their complementary nature in action understanding when compared against video only trained model. This analysis is shown in Figure 10.